OpenStudy (anonymous):

I WILL FAN AND MEDAL Jake's band played songs at three different parties. At the first party they played four-fifths as long as they did at the second party, which was one-half as long as the third party. If they played for 6 minutes at the first party, how many minutes did they play at the third party? Choose one answer. A. 2 minutes B. four and four-fifths minutes C. eight and three-fifths minutes D. fifteen minutes

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (teddyiswatshecallsme):

So to make it all a bit simpler, you represent the first party as 'a', second party as 'b' and third party as 'c'. a=4/5b c = 1/2 b since a=6, 6=4/5(b) b=7.5 c=2(7.5) c=15 So they played 15 minutes at the third party.

OpenStudy (anonymous):

ur amazing and a lifesavor

OpenStudy (teddyiswatshecallsme):

Thank you!